You are getting a discount interest loan of $25,000 for one year. The stated rate of interest is 6%. Calculate the effective interest rate ______%.

  • A. 6.83
  • B. 6.38
  • C. 6.12
  • D. 6.00

You can borrow $75,000 for six months at a stated annual rate of 8%. Calculate the effective annual interest rate _______%.

  • A. 16.64
  • B. 8.50
  • C. 8.16
  • D. 8.00
Answer #1

a. The effective interest rate will be as follows:

= ( 1 + 0.06 / 1 )1 - 1

= 6.00%

So the correct answer is option D i.e. 6.00

b. The effective interest rate will be as follows:

= ( 1 + r / m )m -1

= ( 1 + 0.08 / 2 )2 - 1

= 8.16%

So the correct answer is option C i.e. 8.16%

Feel free to ask in case of any query relating to this question